there are two numbers whose sum is 53. three times the smaller number is equal to 19 more than the large number. What are the numbers?
Answer by mananth(16949)

let smaller number be x
larger number be y
x+y=53.................1
3x=y+19
3x-y=19................2
add equation 1&2
x+y+3x-y=53+19
4x=72
divide by 4
4x/4 = 72/4
x=18
plug the value of x in equation 1
18+y=53
y=53-18
y=35